If in a triangle a=25,b=52,c=63 then the value of tan A/2,tan B/2 is

Grade:12

1 Answers

Shwati
6 years ago
The following are the properties of a triangle.Let a,b,c denote the three sides of the triangle and A,B,C be their opposite angles. We havetanA2=(s−b)(s−c)s(s−a) ...√tan?A2=(s−b)(s−c)s(s−a)tanB2=(s−a)(s−c)s(s−b)... √tan?B2=(s−a)(s−c)s(s−b)tanC2=(s−b)(s−a)s(s−c).... √tan?C2=(s−b)(s−a)s(s−c)where s=a+b+c2